Calculate the pressure needed for the pump of a 3 cm radius fire hose to pump water through 

its nozzle of radius 2 cm at an average speed of 4 m/s when the nozzle is 15 m above the fire 

track. 


1
Expert's answer
2021-12-06T09:40:47-0500
"3^2v=2^2(4)\\v=1.78\frac{m}{s}\\P+0.5(1000)(1.78)^2\\=0.5(1000)(4)^2+(1000)(9.8)(15)\\P=153416\ Pa"